OpenStudy (anonymous):

BEST ANSWER GETS THE MEDAL!!!!1 Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 4)(x − 4) and (x + 4)(x + 4)

OpenStudy (aravindg):

(x + 4)(x − 4) Apply (a+b)(a-b)=a^2-b^2

OpenStudy (aravindg):

(x + 4)(x + 4)=(x + 4)^2

OpenStudy (anonymous):

u can form the sentences now

OpenStudy (anonymous):

THANKS

OpenStudy (aravindg):

yw :)

OpenStudy (anonymous):

HEY I STILL KINDA DONT UNDERSTAND

OpenStudy (aravindg):

caps off please

OpenStudy (anonymous):

can you walk me through how to slove it

OpenStudy (anonymous):

@AravindG